# SciTS: Scientific Time Series Understanding and Generation with LLMs

### v1

@@ 7 Conclusions
We introduce SciTS, a benchmark for scientific time series understanding and generation, spanning 12 disciplines, 7 task types, 43 domain-specific tasks, and over 50k instances, with a wide range of frequencies, lengths, and multivariate dimensions. We benchmark 17 models on SciTS, including text-only LLMs, multimodal LLMs, and unified time series models. Results show that LLMs exhibit stronger adaptability and generalisation compared with models specialised for time series. Two key insights emerge. First, time series should be treated as a dedicated modality: serialising temporal signals as text or rasterising them as images incurs information loss and hinders efficiency and numerical fidelity. Second, while general-purpose LLMs generalise better to unseen domains, pairing them with dedicated temporal processing is crucial to handle complex scientific dynamics. To this end, we propose TimeOmni, a framework that integrates time series into LLMs while remaining compatible with general-purpose LLM training, supporting both understanding and generation. By providing both a comprehensive benchmark and a modelling approach, this work takes a step toward LLMs that can faithfully interrogate dynamical systems, enable discovery, and guide experimental design on real-world scientific temporal data.
@@ Acknowledgments
We thank the PrismaX Team [^4] https://huggingface.co/PrismaX for their support and valuable advice in building SciTS.
@@
Due to resource constraints, we evaluate all models in a zero-shot setting without finetuning. In addition, our preliminary small-scale experiments indicate that enabling the “thinking” mode in closed-source LLMs does not improve performance on time-series tasks. Since it also incurs substantial computational cost, we disable this option for model assessments.
@@
Additional statistics and details of SciTS are provided in Appendix D, with case examples in Appendix F. A detailed description of the assessed models is given in Appendix E, and implementation details of TimeOmni are presented in Appendix C. All datasets, code, and model checkpoints will be made publicly available upon acceptance. A subset of benchmark samples is included in the supplementary material for illustration purposes.
@@ B.1 Time Series Models
Unified Time Series Models. Moirai (Woo et al., 2024) propose the Masked Encoder-based Universal Forecasting Transformer architecture, capable of processing various time series signals. Models ba Building on this, Moirai-MoE (Xu et al., 2025a) and Time-MoE (Shi et al., 2025) integrate Mixture of Experts (MoE) into the Transformer framework, enabling specialized subnetworks to capture heterogeneous temporal patterns through dynamic routing. Concurrently, TimeMixer++ (Wang et al., 2025b) advances the paradigm by capturing task-adaptive temporal patterns through a multi-scale decomposition strategy, simultaneously modelling intricate dependencies in both the time and frequency domains. This dual-domain analysis allows the model to extract features at varying temporal resolutions, enhancing its ability to generalize across tasks with differing periodicities and noise profiles. In contrast to architectural overhauls of large Transformer backbones, SimpleTM (Chen et al., 2025a) adopts a more principled and minimalist design philosophy. It integrates classical signal processing concepts—such as local averaging and trend-seasonal decomposition—directly into a lightweight neural architecture, complemented by a carefully modified attention mechanism. Rather than introducing complex modifications to a pre-existing large-scale model, SimpleTM incrementally incorporates domain-inspired components from scratch, achieving competitive performance with significantly reduced model complexity. Addressing the pervasive challenge of missing data in real-world time series, S4M (Peng et al., 2025) introduces a prototype-based memory mechanism that leverages rich historical patterns stored in a learnable prototype bank. By encoding both observed values and binary masks indicating missingness, the model learns robust and context-aware representations that are resilient to data incompleteness. The integration of explicit missingness indicators with structured memory enables more accurate imputation and forecasting.
@@ Appendix D More Details of SciTS
Astronomy 2048 24 6.4e4 1.2e-5 1.6e4 1
Earth Science 2048 6.0e3 6.0e3 1.0e2 1.0e2 3
@@
Gait freezing classification PHU05 Anomaly detection
@@ D.1 Tasks Description
  • PHU05-Gait freezing classification: Detecting sudden “freezing” episodes (temporary movement cessation) during walking in patients with Parkinson’s disease, supporting rehabilitation assessment and intervention. The input to this task consists of single-channel time series data and text, and the output is text.
@@
Task ID Discipline Input TS Generation TS Frequency Channel Data Source
lLength Range Length Range Range(Hz) Range
ASU01 Astronomy 6.4e4 / 1.6e4 1 Simulation
ASG02 6.4e4 1 1.6e4 1 Simulation
@@
EAU01 Earth Science 6.0e3 / 1.0e2 3 STEAD (Mousavi et al., 2019)
EAG02 6.0e3 2 1.0e2 3 STEAD (Mousavi et al., 2019)
@@
MEU01 Meteorology 16 / Unknown 1 TS MQA (Kong et al., 2025)
MEU02 24 / 2.8e-4 5 TIMECAP (Lee et al., 2025)
@@
NEU05 1.0e3 / 2.5e2 58 WBCIC SHU (Yang et al., 2025b)
NEU06 3.0e3 / 1.0e2 1 Sleep (Kemp et al., 2000)
@@
PHU04 64 / 5.0e2 1 TS MQA (Kong et al., 2025)
PHU05 9 / 64 1 TS MQA (Kong et al., 2025)
PHU06 30 / 20 1 TS MQA (Kong et al., 2025)
URG01 Urbanism 24 24 2.8e-4 1 NewsForecast (Wang et al., 2024b)
@@
MFU03 1.6e5 / 1.6e4 1 MIMII Due (Tanabe et al., 2021)
RAU01 Radar $3\sim$3.2e2 / 3.2e6 2 RadSeg (Huang et al., 2024)
RAU02 1.3e2 / 1.0e7 2 RadarComm (Jagannath & Jagannath, 2021)
MAG01 Math 1.0e2 20 1.0e2 3 Chaotic (Gilpin, 2023)
@@ D.2 Dataset Description
GWOSC GW Event. Since the number of observed gravitational events is quite limited. We constructed GWOSC GW Event dataset by simulation, following prior work in astronomy LABEL:dax2021real,dax2023neural,dax2025real. To construct synthetic data that closely match real observations, we generated gravitational-wave templates via numerical relativity simulations and injected them into strain data downloaded from the GWOSC platform [^5] https://gwosc.org/eventapi/html/. Unlike the traditional approach of generating noise from a fixed power spectral density (PSD), this method leverages real detector noise segments to enrich noise diversity. Specifically, we applied data-quality and injection masks to exclude low-quality or contaminated intervals, and extracted 4-second noise segments (sampled at 16 kHz) starting 16 s after each event to form a noise pool. A template bank was built through a grid search over typical binary black hole mass ranges. For each template, two noise segments were randomly drawn and scaled to different matched-filter SNRs by adjusting the amplitude ratio. To mitigate truncation artifacts, windowing was applied during waveform-noise superposition, and the final data were whitened to produce realistic and training-ready synthetic datasets.

### v2

@@ 7 Conclusions
We introduce SciTS, a benchmark for scientific time series understanding and generation, spanning 12 disciplines, 7 task types, 43 domain-specific tasks, and over 50k instances, with a wide range of frequencies, lengths, and multivariate dimensions. We benchmark 17 models on SciTS, including text-only LLMs, multimodal LLMs, and unified time series models. Two key insights emerge. First, LLMs exhibit stronger adaptability and generalisation compared with models specialised for time series. Second, while general-purpose LLMs generalise better to unseen domains, pairing them with dedicated temporal processing is crucial to handle complex scientific dynamics. Serialising temporal signals as text or rasterising them as images incurs information loss and hinders efficiency and numerical fidelity. To this end, we provide TimeOmni, a demonstrated framework that integrates time series into LLMs while remaining compatible with general-purpose LLM training. By providing a comprehensive benchmark and a preliminary approach, this work takes a step toward LLMs that can faithfully interrogate dynamical systems, enable discovery, and guide experimental design on real-world scientific temporal data.
@@ Acknowledgments
Supported by Shanghai Artificial Intelligence Laboratory. We thank the PrismaX Team [^4] https://huggingface.co/PrismaX for their support and valuable advice in building SciTS.
@@
Due to resource constraints, we evaluate all baselines in a zero-shot setting without finetuning. More explanation of the evaluation setting can be found in Appendix H. In addition, our preliminary small-scale experiments indicate that enabling the “thinking” mode in closed-source LLMs does not improve performance on time-series tasks. Since it also incurs substantial computational cost, we disable this option for model assessments.
@@
Additional statistics and details of SciTS are provided in Appendix D, with case examples in Appendix I. A detailed description of the assessed models is given in Appendix E, and implementation details of TimeOmni are presented in Appendix C. The SciTS benchmark is available at https://huggingface.co/datasets/OpenTSLab/SciTS. The TimeOmni framework is available at https://github.com/OpenTSLab/TimeOmni.
@@ B.1 Time Series Models
Unified Time Series Models. Moirai (Woo et al., 2024) propose the Masked Encoder-based Universal Forecasting Transformer architecture, capable of processing various time series signals. Models building on this, Moirai-MoE (Xu et al., 2025a) and Time-MoE (Shi et al., 2025) integrate Mixture of Experts (MoE) into the Transformer framework, enabling specialized subnetworks to capture heterogeneous temporal patterns through dynamic routing. Concurrently, TimeMixer++ (Wang et al., 2025b) advances the paradigm by capturing task-adaptive temporal patterns through a multi-scale decomposition strategy, simultaneously modelling intricate dependencies in both the time and frequency domains. This dual-domain analysis allows the model to extract features at varying temporal resolutions, enhancing its ability to generalize across tasks with differing periodicities and noise profiles. In contrast to architectural overhauls of large Transformer backbones, SimpleTM (Chen et al., 2025a) adopts a more principled and minimalist design philosophy. It integrates classical signal processing concepts—such as local averaging and trend-seasonal decomposition—directly into a lightweight neural architecture, complemented by a carefully modified attention mechanism. Rather than introducing complex modifications to a pre-existing large-scale model, SimpleTM incrementally incorporates domain-inspired components from scratch, achieving competitive performance with significantly reduced model complexity. Addressing the pervasive challenge of missing data in real-world time series, S4M (Peng et al., 2025) introduces a prototype-based memory mechanism that leverages rich historical patterns stored in a learnable prototype bank. By encoding both observed values and binary masks indicating missingness, the model learns robust and context-aware representations that are resilient to data incompleteness. The integration of explicit missingness indicators with structured memory enables more accurate imputation and forecasting.
@@ Appendix D More Details of SciTS
Astronomy 2583 24 6.4e4 1.2e-5 1.6e4 1
Earth Science 3080 6.0e3 6.0e3 1.0e2 1.0e2 3
@@
Gait freezing anomaly detection PHU05 Anomaly detection
@@ D.1 Tasks Description
  • PHU05-Gait freezing anomaly detection: Detecting sudden “freezing” episodes (temporary movement cessation) during walking in patients with Parkinson’s disease, supporting rehabilitation assessment and intervention. The input to this task consists of single-channel time series data and text, and the output is text.
@@
Discipline Task ID # Instances Input TS Generation TS Frequency Channel Data Source
Length Range Length Range Range(Hz) Range
Astronomy ASU01 1024 6.4e4 / 1.6e4 1 Simulation
ASG02 535 6.4e4 1 1.6e4 1 Simulation
@@
Earth Science EAU01 2048 6.0e3 / 1.0e2 3 STEAD (Mousavi et al., 2019)
EAG02 1032 6.0e3 2 1.0e2 3 STEAD (Mousavi et al., 2019)
@@
Meteorology MEU01 1256 16 / Unknown 1 TS MQA (Kong et al., 2025)
MEU02 1131 24 / 2.8e-4 5 TIMECAP (Lee et al., 2025)
@@
NEU05 3600 1.0e3 / 2.5e2 58 WBCIC SHU (Yang et al., 2025b)
NEU06 1024 3.0e3 / 1.0e2 1 Sleep (Kemp et al., 2000)
@@
PHU04 1006 64 / 5.0e2 1 TS MQA (Kong et al., 2025)
PHU05 1882 9 / 64 1 TS MQA (Kong et al., 2025)
PHU06 $18$ 30 / 20 1 TS MQA (Kong et al., 2025)
Urbanism URG01 43 24 24 2.8e-4 1 NewsForecast (Wang et al., 2024b)
@@
MFU03 3600 1.6e5 / 1.6e4 1 MIMII Due (Tanabe et al., 2021)
Radar RAU01 963 $3\sim$3.2e2 / 3.2e6 2 RadSeg (Huang et al., 2024)
RAU02 1280 1.3e2 / 1.0e7 2 RadarComm (Jagannath & Jagannath, 2021)
Math MAG01 1856 1.0e2 20 1.0e2 3 Chaotic (Gilpin, 2023)
@@ D.2 Dataset Description
GWOSC GW Event. Since the number of observed gravitational events is quite limited. We constructed GWOSC GW Event dataset by simulation, following prior work in astronomy (Dax et al., 2021; 2023; 2025). To construct synthetic data that closely match real observations, we generated gravitational-wave templates via numerical relativity simulations and injected them into strain data downloaded from the GWOSC platform [^5] https://gwosc.org/eventapi/html/. Unlike the traditional approach of generating noise from a fixed power spectral density (PSD), this method leverages real detector noise segments to enrich noise diversity. Specifically, we applied data-quality and injection masks to exclude low-quality or contaminated intervals, and extracted 4-second noise segments (sampled at 16 kHz) starting 16 s after each event to form a noise pool. A template bank was built through a grid search over typical binary black hole mass ranges. For each template, two noise segments were randomly drawn and scaled to different matched-filter SNRs by adjusting the amplitude ratio. To mitigate truncation artifacts, windowing was applied during waveform-noise superposition, and the final data were whitened to produce realistic and training-ready synthetic datasets.